OpenStudy (anonymous):

Tiya flipped a coin 40 times. The coin landed heads up 16 times and tails up 24 times. Part A: Based on the results, what is the experimental probability of the coin landing heads up? Show your work. (5 points) Part B: What is the theoretical probability of the coin landing heads up? Show your work. (5 points) @

OpenStudy (anonymous):

@Michele_Laino help im dying of help????

OpenStudy (anonymous):

@Michele_Laino help????????????????????

OpenStudy (anonymous):

@Michele_Laino are you goign to help me ??

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

part A: the requested probability, is given by definition, by the subsequent formula: p=favorable cases/possible cases=16/40=...?

OpenStudy (anonymous):

so i divide @Michele_Laino

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

what about part b

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

part B: we have favorable cases = 1, since we have only one head face of our coin and possible cases = 2, since our coin has 2 faces, so probability=favorable cases/possible cases= 1/2=...?

OpenStudy (anonymous):

50 percent @Michele_Laino

OpenStudy (anonymous):

2 more questions @Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Maria spins a penny 100 times and it lands head side up 62 times. Explain why Maria's experimental probability may be different from the theoretical probability of spinning a coin.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the experimental probability, is: p= favorable cases/possible cases=62/100=...?

OpenStudy (anonymous):

Explain why Maria's experimental probability may be different from the theoretical probability of spinning a coin. @Michele_Laino

OpenStudy (michele_laino):

yes! I understand, nevertheless in order to answer to your question, it is necessary to compute the experimental probability. So what is: p=favorable cases/possible cases=62/100=...?

OpenStudy (anonymous):

62%

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now the theoretical probability is as in the previous exercise, namely: p=favorable cases/possible cases=1/2=...?

OpenStudy (anonymous):

50 % @Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

that's the reason, since 62% is different from 50%

OpenStudy (michele_laino):

is it ok?

OpenStudy (anonymous):

Huh @Michele_Laino

OpenStudy (michele_laino):

since we have computed the experimental probability and it is 62% after that we have computed the theoretical probability, and it is 50% Now we can see that those probabilities are different each other

OpenStudy (anonymous):

but that dosent explain @Michele_Laino

OpenStudy (michele_laino):

why?

OpenStudy (anonymous):

ok next question ?

OpenStudy (anonymous):

Jake tossed a paper cup 50 times and recorded how it landed. The table shows the results: Position Open Side Up Closed Side Up Landing on Side Number of Times Landed in Position 1 5 44 Based on the table, determine the experimental probability of each outcome (landing open side up, landing closed side up, and landing on its side). Show your work.

OpenStudy (michele_laino):

a possible reason for which those probability are different can be this: our coin is not a "honest" coin, namely our coin is a rigged coin

OpenStudy (anonymous):

@Michele_Laino Jake tossed a paper cup 50 times and recorded how it landed. The table shows the results: Position Open Side Up Closed Side Up Landing on Side Number of Times Landed in Position 1 5 44 Based on the table, determine the experimental probability of each outcome (landing open side up, landing closed side up, and landing on its side). Show your work.

OpenStudy (michele_laino):

oops.. those probabilities*

OpenStudy (michele_laino):

I'm trying to solve your exercise, please wait

OpenStudy (michele_laino):

here we have possible cases = 1+ 5+44=50

OpenStudy (michele_laino):

probability for "open side up" event: favorable cases/possible cases = 1/50=...?

OpenStudy (anonymous):

2%

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

probability for "closed side up": p= favorable cases/possible cases= 5/50=...?

OpenStudy (anonymous):

10%

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

finally, probability for "landing on side" event: p=favorable cases/possible cases=44/50=...?

OpenStudy (anonymous):

88% @Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

one last question

OpenStudy (anonymous):

Brenda throws a dart at this square-shaped target: A square is shown with sides labeled 11. A shaded circle is shown in the center of the square. The diameter of the circle is 2. Part A: Is the probability of hitting the black circle inside the target closer to 0 or 1? Explain your answer. (5 points) Part B: Is the probability of hitting the white portion of the target closer to 0 or 1? Explain your answer. (5 points)

OpenStudy (anonymous):

i just need help with part b

OpenStudy (michele_laino):

what is the side of the square?

OpenStudy (michele_laino):

is it 11?

OpenStudy (anonymous):

yes @Michele_Laino

OpenStudy (michele_laino):

 the area of white portion is: 11*11-3.14*1*1=117.86 so the requested probability, is: p=117.86/121=...?

OpenStudy (anonymous):

0/974

OpenStudy (anonymous):

0.974 @Michele_Laino

OpenStudy (michele_laino):

and such probability is closer to 1

OpenStudy (anonymous):

how

OpenStudy (anonymous):

@michele_laino

OpenStudy (michele_laino):

since 0.974 is close to 1